# VL-JEPA: Joint Embedding Predictive Architecture for Vision-language Delong Chen1,2,* Mustafa Shukor1,3,* Théo Moutakanni1,* Willy Chung1,3,* Jade Yu1, Tejaswi Kasarla1, Allen Bolourchi1, Yann LeCun1,4, Pascale Fung1,2

### v1

@@
Delong Chen1,2,* Mustafa Shukor1,3,* Théo Moutakanni1,* Willy Chung1,3,*
Jade Yu1, Tejaswi Kasarla1, Allen Bolourchi1, Yann LeCun1,4, Pascale Fung1,2
@@ 2 Methodology
Fig. 2 illustrates how we instantiate the VL-JEPA architecture in this paper. For the X-Encoder, we chose V-JEPA 2 (Assran et al., 2025), a Vision Transformer that outputs a sequence of visual tokens, which are then projected and fed into the Predictor initialized using Llama 3 Transformer layers. Query conditioning is achieved by tokenizing and embedding the textual query and feeding the resulting textual token embeddings into the Predictor along with the visual embeddings. The outputs of the Llama 3 Transformer layers are pooled and projected into the target embedding space produced by the Y-Encoder, which is initialized by EmbeddingGemma-300M (Vera et al., 2025). We provide more technical details in §LABEL:sec:implementation_details.
@@ 3.2 Two-stage Training
Large-scale Pretraining. VL-JEPA is trained with two stages. The first query-free pretraining stage aims to establish robust vision-language alignment using massive caption data. We use PLM-Image-Auto (Cho et al., 2025), Datacomp (Gadre et al., 2023) and YFCC-100M (Thomee et al., 2016) for image-text data. For video-text data, we include PLM-Video-Auto (Cho et al., 2025), Ego4D atomic action descriptions (Grauman et al., 2022), and an internal dataset Action100M consisting captions generated on HowTo100M videos (Chen et al., 2025b).
We first do image-only training on Datacomp and YFCC-100M with only 1 frame per visual input, which allows us to use a large batch size of 24k. After 100k iterations, the model has seen 2B samples and achieved 61.6% ImageNet zero-shot accuracy (without prompt ensembling). Then, we continue with joint image-video pretraining with 16 frames per input. The pretraining takes 2 weeks using 24 nodes with 8NVIDIA H200 GPUs each. We adopt a constant learning rate of to facilitate extended training. We call the resulting model VL-JEPA and measure zero-shot classification and retreival performance with this model.
@@
We train the model for 35k steps with a batch size of 6k (2 days with 24 nodes), with cosine learning rate annealing applied to improve convergence. Since excessive human labelled data is included in this SFT data mixture, we no longer emphasize zero-shot evaluation for the resulting VL-JEPA from this stage. Instead, we evaluate VQA capabilities and compare it with state-of-the-art specialist models.
@@
SSv2 |
EK100 |
EgoExo4D |
Kinetics-400 |
COIN (SR) |
COIN (TR) |
CrossTask (SR) |
CrossTask (TR) |
Average |
MSR-VTT |
ActivityNet |
DiDeMo |
MSVD |
YouCook2 |
PVD-Bench |
Dream-1k |
VDC-1k |
| RN50 | 75M | 12.8B | 21.8 | 2.1 | 1.5 | 1.9 | 41.4 | 8.6 | 39.0 | 10.9 | 68.7 | 28.3 | 28.7 | 17.7 | 24.7 | 29.7 | 5.1 | 27.6 | 47.2 | 46.0 | |||
| ViT-B | 124M | 12.8B | 25.3 | 3.1 | 1.3 | 2.4 | 49.5 | 11.2 | 47.3 | 16.2 | 71.5 | 29.3 | 31.0 | 19.5 | 25.7 | 34.0 | 6.1 | 27.0 | 48.5 | 42.9 | |||
| CLIP | ViT-L | 389M | 12.8B | ✓ | ✓ | 30.9 | 3.8 | 3.7 | 3.6 | 58.3 | 14.7 | 63.5 | 20.8 | 78.5 | 35.3 | 35.9 | 23.4 | 30.7 | 41.9 | 7.9 | 36.7 | 56.8 | 49.3 |
| ViT-B | 375M | 40B | 33.9 | 5.2 | 2.3 | 4.9 | 57.8 | 20.6 | 69.9 | 27.7 | 82.9 | 39.6 | 40.2 | 25.0 | 32.1 | 48.6 | 13.8 | 52.1 | 60.9 | 43.7 | |||
| ViT-L | 882M | 40B | 38.7 | 5.9 | 4.5 | 7.0 | 63.6 | 24.2 | 78.5 | 35.1 | 90.8 | 45.4 | 41.6 | 32.7 | 35.1 | 53.5 | 19.0 | 59.2 | 71.6 | 50.9 | |||
| SigLIP2 | ViT-g | 1.9B | 40B | ✓ | ✓ | 39.9 | 6.1 | 6.1 | 6.4 | 68.0 | 26.0 | 80.4 | 35.1 | 90.8 | 47.5 | 43.4 | 33.9 | 38.9 | 56.0 | 22.2 | 60.4 | 73.0 | 52.5 |
| ViT-B | 448M | 58B | 37.3 | 5.8 | 3.3 | 6.3 | 65.4 | 21.5 | 77.1 | 26.9 | 91.8 | 44.9 | 46.5 | 35.4 | 35.3 | 49.1 | 15.2 | 59.8 | 68.7 | 49.2 | |||
| ViT-L | 671M | 58B | 42.8 | 9.3 | 6.0 | 10.9 | 73.4 | 27.1 | 83.3 | 37.5 | 95.3 | 50.2 | 48.9 | 41.7 | 40.8 | 56.2 | 22.5 | 64.7 | 75.9 | 51.0 | |||
| PE-Core | ViT-G | 2.3B | 86B | ✓ | ✓ | 44.6 | 9.0 | 6.4 | 13.0 | 76.4 | 29.0 | 86.0 | 40.3 | 97.2 | 58.1 | 51.6 | 49.1 | 44.5 | 58.7 | 26.0 | 77.0 | 89.2 | 68.5 |
| VL-JEPA | ViT-L | 1.6B | 2.0B | ✓ | ✓ | 46.4 | 16.1 | 13.3 | 21.1 | 57.8 | 39.8 | 74.4 | 60.5 | 88.0 | 58.4 | 37.6 | 55.4 | 49.2 | 47.9 | 23.1 | 78.2 | 88.8 | 87.2 |
| VL-JEPA | ViT-L | 1.6B | 2.5B | ✗ | ✓ | 70.7 | 68.2 | 38.8 | 59.5 | 81.4 | 60.3 | 86.8 | 77.1 | 93.0 | 59.5 | 43.7 | 53.8 | 46.2 | 49.1 | 28.8 | 81.1 | 86.4 | 86.7 |
| SoTA (including specialist models) | ✗ | ✗ | - | 77.5 | 56.4 | 47.8 | 92.1 | 67.3 | 95.3 | 64.5 | 96.0 | - | 62.8 | 74.1 | 74.2 | 61.4 | 28.9 | 77.0 | 89.2 | 68.5 | |||
@@ 4.1 Classification and Retrieval
We begin by evaluating VL-JEPA’s classification and retrieval performance in §4.1, and benchmark VL-JEPA on VQA datasets in §4.2. We demonstrate application of VL-JEPA for understanding the relationship between world state changes and action concepts (i.e., inverse dynamics) in §4.3. In §4.4, we demonstrate the advantage of embedding prediction by comparing it with a token-predictive VLM baseline under a strictly controlled setting. In §4.5, we evaluate the effectiveness of VL-JEPA’s selective decoding, and show that it reduces decoding cost while maintaining the performance. Next, we analyze VL-JEPA’s Y-Encoder in §4.6.
Evaluation Setup. We evaluate VL-JEPA following the CLIP-style evaluation protocol (see Fig.2 and §2 “Multi-tasking”). We assess VL-JEPA on a broad suite of benchmarks, including 8 classification datasets and 8 retrieval datasets. For zero-shot evaluation, we compare against generalist foundation models CLIP (Radford et al., 2021), SigLIP2 (Tschannen et al., 2025), and Perception Encoder (PE-Core)(Bolya et al., 2025). We additionally report reference numbers from specialist models that are individually optimized for each benchmark (summarized in AppendixLABEL:sec:app.cls-ret).
Results. Table 1 summarizes the results. In the strict zero-shot setting, VL-JEPA achieves higher average accuracy (46.4 vs 44.6) across the 8 classification datasets and higher average recall@1 (58.4 vs 58.1) across the 8 retrieval datasets than the best baseline PE-Core-G. Per-dataset scores show that VL-JEPA is particularly strong on motion-centric benchmarks (SSv2, EK-100, EgoExo4D, and step recognition on COIN and CrossTask), while relatively weaker on appearance-centric benchmarks (Kinetics-400 and task recognition on COIN and CrossTask). This is due to VL-JEPA has seen substantially fewer vision-language pairs (only 2B in comparison with PE-Core-G’s 86B). After supervised finetuning, VL-JEPA improves significantly upon VL-JEPA since the model has seen in-domain training data. As a single generalist model, the performance of VL-JEPA is approaching specialist models optimized individually for each dataset.
@@
| VL-JEPA (1.6B) | 60.8 | VL-JEPA (1.6B) | 67.4 | VL-JEPA (1.6B) | 84.2 | VL-JEPA (1.6B) | 82.2 | |||
@@ 5 Related Works
JEPA Models. JEPA model learns by predicting the representation of a target input from the representation of a context input . Early instantiations include I-JEPA for image encoding (Assran et al., 2023) and V-JEPA for video encoding (Bardes et al., 2023), which demonstrated the effectiveness of this objective over pixel reconstruction approach in their respective modality. Recent JEPA work falls into two categories. One category of work emphasizes better unimodal representation learning (Assran et al., 2023; Bardes et al., 2023; Fei et al., 2023) or cross-modal alignment (Lei et al., 2025; Jose et al., 2025). The other direction targets world modeling, where pretrained encoders are frozen and action-conditioned predictors are trained for conditional prediction of state representations (Zhou et al., 2025; Baldassarre et al., 2025; Assran et al., 2025). This has shown good results but remains limited to narrow domains like mazes or robotic pick-and-place. Our proposed VL-JEPA is the first designed for general-purpose vision–language tasks. It performs conditional latent prediction over vision and text, and preserves efficiency while enabling flexible, multitask architecture.
@@ Acknowledgments
We would like to thank Yejin Bang, Adrien Bardes, Loïc Barrault, Lucas Beyer, Quentin Garrido, João Maria Janeiro, Yifu Qiu, Koustuv Sinha, Basile Terver, and François Yvon for providing valuable feedback and support to this work.

### v2

@@
Delong Chen1,2*, Mustafa Shukor1,3*, Théo Moutakanni1*, Willy Chung1,3*,
Jade Yu1, Tejaswi Kasarla1, Yejin Bang1, Allen Bolourchi1, Yann LeCun1,4, Pascale Fung1,2
@@ 2 Methodology
Fig. 2 illustrates how we instantiate the VL-JEPA architecture in this paper. For the X-Encoder, we chose V-JEPA 2 (Assran et al., 2025), a Vision Transformer that outputs a sequence of visual tokens, which are then projected and fed into the Predictor initialized using Llama 3 Transformer layers. Query conditioning is achieved by tokenizing and embedding the textual query and feeding the resulting textual token embeddings into the Predictor along with the visual embeddings. The outputs of the Llama 3 Transformer layers are pooled and projected into the target embedding space produced by the Y-Encoder, which is initialized by EmbeddingGemma-300M (Vera et al., 2025). We provide more technical details in §3.
@@ 3.2 Two-stage Training
Large-scale Pretraining. VL-JEPA is trained with two stages. The first query-free pretraining stage aims to establish robust vision-language alignment using massive caption data. We use Datacomp (Gadre et al., 2023) and YFCC-100M (Thomee et al., 2016) for image-text data. For video-text data, we use Action100M (Chen et al., 2026), which consists action description and video captions generated on HowTo100M videos (Chen et al., 2025b).
We first do image-only training on Datacomp and YFCC-100M with only 1 frame per visual input, which allows us to use a large batch size of 24k. After 100k iterations, the model has seen 2B samples and achieved 61.6% ImageNet zero-shot accuracy (without prompt ensembling). Then, we continue with video pretraining with 8 frames per input for 60k iterations and 32 frames for 10k iterations in the end. The pretraining takes 4 weeks using 24 nodes with 8NVIDIA H200 GPUs each. We adopt a constant learning rate of to facilitate extended training. We refer the readers to the Action100M paper (Chen et al., 2026) for more details. We call the resulting model VL-JEPA and measure zero-shot classification and retrieval performance with this model.
@@
We train the model for 83k steps with a batch size of 3,072 (2.5s days with 24 nodes), with cosine learning rate annealing applied to improve convergence. Since excessive human labeled data is included in this SFT data mixture, we no longer emphasize zero-shot evaluation for the resulting VL-JEPA from this stage. Instead, we evaluate VQA capabilities and compare it with state-of-the-art specialist models.
@@
SSv2
(Goyal et al., 2017)
EK100
(Damen et al., 2022)
EgoExo4D
(Grauman et al., 2024)
Kinetics-400
(Kay et al., 2017)
COIN (SR)
(Tang et al., 2019)
COIN (TR)
(Tang et al., 2019)
CrossTask (SR)
(Zhukov et al., 2019)
CrossTask (TR)
(Zhukov et al., 2019)
Average
MSR-VTT
(Xu et al., 2016)
ActivityNet
(Caba Heilbron et al., 2015)
DiDeMo
(Anne Hendricks et al., 2017)
MSVD
(Chen and Dolan, 2011)
YouCook2
(Zhou et al., 2018)
PVD-Bench
(Bolya et al., 2025)
Dream-1k
(Wang et al., 2024)
VDC-1k
(Chai et al., 2024)
| RN50 | 75M | 12.8B | 21.8 | 2.1 | 1.5 | 2.1 | 41.4 | 8.6 | 39.0 | 10.9 | 68.7 | 28.3 | 28.7 | 17.7 | 24.7 | 29.7 | 5.1 | 27.6 | 47.2 | 46.0 | |||
| ViT-B | 124M | 12.8B | 25.4 | 3.1 | 1.3 | 2.8 | 49.5 | 11.2 | 47.3 | 16.2 | 71.5 | 29.3 | 31.0 | 19.5 | 25.7 | 34.0 | 6.1 | 27.0 | 48.5 | 42.9 | |||
| CLIP | ViT-L | 389M | 12.8B | ✓ | ✓ | 30.7 | 3.8 | 3.7 | 2.6 | 58.3 | 14.7 | 63.5 | 20.8 | 78.5 | 35.3 | 35.9 | 23.4 | 30.7 | 41.9 | 7.9 | 36.7 | 56.8 | 49.3 |
| ViT-B | 375M | 40B | 33.9 | 5.2 | 2.3 | 4.5 | 57.8 | 20.6 | 69.9 | 27.7 | 82.9 | 39.6 | 40.2 | 25.0 | 32.1 | 48.6 | 13.8 | 52.1 | 60.9 | 43.7 | |||
| ViT-L | 882M | 40B | 38.6 | 5.9 | 4.5 | 6.4 | 63.6 | 24.2 | 78.5 | 35.1 | 90.8 | 45.4 | 41.6 | 32.7 | 35.1 | 53.5 | 19.0 | 59.2 | 71.6 | 50.9 | |||
| SigLIP2 | ViT-g | 1.9B | 40B | ✓ | ✓ | 39.8 | 6.1 | 6.1 | 5.6 | 68.0 | 26.0 | 80.4 | 35.1 | 90.8 | 47.5 | 43.4 | 33.9 | 38.9 | 56.0 | 22.2 | 60.4 | 73.0 | 52.5 |
| ViT-B | 448M | 58B | 37.2 | 5.8 | 3.3 | 6.0 | 65.4 | 21.5 | 77.1 | 26.9 | 91.8 | 44.9 | 46.5 | 35.4 | 35.3 | 49.1 | 15.2 | 59.8 | 68.7 | 49.2 | |||
| ViT-L | 671M | 58B | 42.9 | 9.3 | 6.0 | 11.6 | 73.4 | 27.1 | 83.3 | 37.5 | 95.3 | 50.2 | 48.9 | 41.7 | 40.8 | 56.2 | 22.5 | 64.7 | 75.9 | 51.0 | |||
| PE-Core | ViT-G | 2.3B | 86B | ✓ | ✓ | 44.7 | 9.0 | 6.4 | 13.6 | 76.4 | 29.0 | 86.0 | 40.3 | 97.2 | 58.1 | 51.6 | 49.1 | 44.5 | 58.7 | 26.0 | 77.0 | 89.2 | 68.5 |
| VL-JEPA | ViT-L | 1.6B | 3.3B | ✓ | ✓ | 52.5 | 19.3 | 21.8 | 33.2 | 64.8 | 47.4 | 79.4 | 64.5 | 89.6 | 63.7 | 40.0 | 64.9 | 50.0 | 49.0 | 40.4 | 83.1 | 93.3 | 88.8 |
| VL-JEPA | ViT-L | 1.6B | 3.6B | ✗ | ✓ | 75.4 | 73.2 | 44.6 | 68.1 | 84.8 | 66.4 | 90.3 | 79.8 | 96.2 | 63.8 | 46.2 | 62.4 | 52.3 | 52.3 | 37.8 | 82.6 | 89.3 | 87.7 |
| Previous SoTA (including specialist models) | ✗ | ✗ | - | 77.5 | 56.4 | 47.8 | 92.1 | 67.3 | 95.3 | 64.5 | 96.0 | - | 62.8 | 74.1 | 74.2 | 61.4 | 28.9 | 77.0 | 89.2 | 68.5 | |||
@@ 4 Experiments
We begin by evaluating VL-JEPA’s classification and retrieval performance in §4.1, and benchmark VL-JEPA on VQA datasets in §4.2. We demonstrate application of VL-JEPA for understanding the relationship between world state changes and action concepts (i.e., inverse dynamics) in §4.3. In §4.5, we demonstrate the advantage of embedding prediction by comparing it with a token-predictive VLM baseline under a strictly controlled setting. In §4.6, we evaluate the effectiveness of VL-JEPA’s selective decoding, and show that it reduces decoding cost while maintaining the performance. Next, we analyze VL-JEPA’s Y-Encoder in §4.7. Finally, we present ablation studies in §4.8.
@@ 4.1 Classification and Retrieval
Evaluation Setup. We evaluate VL-JEPA following the CLIP-style evaluation protocol (see Fig.2 and §2 “Multi-tasking”). We assess VL-JEPA on a broad suite of benchmarks, including 8 classification datasets and 8 retrieval datasets. For zero-shot evaluation, we compare against generalist foundation models CLIP (Radford et al., 2021), SigLIP2 (Tschannen et al., 2025), and Perception Encoder (PE-Core)(Bolya et al., 2025). We additionally report reference numbers from specialist models that are individually optimized for each benchmark.
Results. Table 1 summarizes the results. In the strict zero-shot setting, VL-JEPA achieves higher average accuracy (52.5 vs 44.7) across the 8 classification datasets and higher average recall@1 (63.7 vs 58.1) across the 8 retrieval datasets than the best baseline PE-Core-G. Per-dataset scores show that VL-JEPA is particularly strong on motion-centric benchmarks (SSv2, EK-100, EgoExo4D, and step recognition on COIN and CrossTask), while relatively weaker on appearance-centric benchmarks (Kinetics-400 and task recognition on COIN and CrossTask). This is due to VL-JEPA has seen substantially fewer vision-language pairs (only 3.6B in comparison with PE-Core-G’s 86B). After supervised finetuning, VL-JEPA improves significantly upon VL-JEPA since the model has seen in-domain training data. As a single generalist model, the performance of VL-JEPA is approaching specialist models optimized individually for each dataset.
@@
| VL-JEPA (1.6B) | 61.5 | VL-JEPA (1.6B) | 69.9 | VL-JEPA (1.6B) | 85.7 | VL-JEPA (1.6B) | 86.3 | |||
@@ 4.3 WorldPrediction-WM
4.4 Action Anticipation
| Model | Vision Encoder | ||||
| V-JEPA2 | ViT-g-384px | 39.7 | 28.6 | 19.3 | 8.2 |
| V-JEPA2 | ViT-L-256px | 32.7 | 23.4 | 15.9 | 7.1 |
| VL-JEPA | ViT-L-256px | 34.2 | 26.0 | 19.4 | 11.7 |
| Model |
49.1 | 49.7 | 53.6 | 56.2 |
Evauation Setup. We assess the action anticipation (i.e., forecasting future actions) capabilities of fine-tuned VL-JEPA on two benchmarks: action anticipation with EPIC-KITCHENS-100 (Damen et al., 2022), and COIN (Tang et al., 2019) next step forecasting. In this task, given a window of context video, the model is required to predict the next action that will occur after a specified anticipation time—the temporal gap between the end of the context segment and the onset of the subsequent action. VL-JEPA is fine-tuned and we compare our results against the state-of-the-art baselines.
Results As compated in Table VL-JEPA demonstrates strong performance in both benchmarks (Table 4 and Table 5). For the Epic-Kitchens-100, particularly at the standard 1-second anticipation interval, where it achieves a Recall@5 of 34.18 – surpassing the previous V-JEPA 2 that uses the same ViT-L-256px encoder by 1.48 points. While the larger V-JEPA 2 ViT-g-384px model attains the highest score at 1 second, VL-JEPA remains competitive and shows notable advantages at longer anticipation intervals. On the COIN next step forecasting task, as shown in Tab. 5, VL-JEPA achieves 56.2%, outperform strong VLM baselines VideoLLM-online (Chen et al., 2024c), VideoLLM-MoD (Wu et al., 2024), ProVideLLM (Chatterjee et al., 2025). Overall, these findings demonstrated that VL-JEPA can be finetuned to handle semantic uncertain prediction tasks.
@@ 5 Related Works
JEPA Models. JEPA model learns by predicting the representation of a target input from the representation of a context input . Early instantiations include I-JEPA for image encoding (Assran et al., 2023) and V-JEPA for video encoding (Bardes et al., 2023), which demonstrated the effectiveness of this objective over pixel reconstruction approach in their respective modality. Recent JEPA work falls into two categories. One category of work emphasizes better unimodal representation learning (Assran et al., 2023; Bardes et al., 2023; Fei et al., 2023) or cross-modal alignment (Lei et al., 2025; Jose et al., 2025). The other direction targets world modeling, where pretrained encoders are frozen and action-conditioned predictors are trained for conditional prediction of state representations (Zhou et al., 2025; Baldassarre et al., 2025; Assran et al., 2025; Terver et al., 2025). This has shown good results but remains limited to narrow domains like mazes or robotic pick-and-place. Our proposed VL-JEPA is the first designed for general-purpose vision–language tasks. It performs conditional latent prediction over vision and text, and preserves efficiency while enabling flexible, multitask architecture.
@@ Acknowledgments
We would like to thank Loïc Barrault, Lucas Beyer, Quentin Garrido, João Maria Janeiro, Yifu Qiu, Koustuv Sinha, Basile Terver, and François Yvon for providing valuable feedback and support to this work. We thank Adrien Bardes for detailed V-JEPA 2 baseline performance on EK-100 action anticipation. We acknowledge Alejandro Castillejo Muñoz for his efforts on real-time demonstration and qualitative evaluation of VL-JEPA.